OpenStudy (anonymous):

cos squared x + 2cosx + 1 = 0

OpenStudy (anonymous):

cos^2 x + cos x + cos x +1 = 0 cos x (cosx +1) + 1 (cos x +1)= 0 (cos x+1)(cos x +1 ) = 0

OpenStudy (anonymous):

i think you do it now

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i have another

OpenStudy (anonymous):

what is your answer

OpenStudy (anonymous):

\[\cos ^{2}x = 3\] is that correct?

OpenStudy (anonymous):

cos x + 1=0 cos x = -1 cos x = cos 2 Pi

OpenStudy (anonymous):

oh ok tnx!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

@WATTAGWAN are ypu still answering? .. tnx

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

ok